# CurvFed: Curvature-Aligned Federated Learning for Fairness without Demographics

### v6

@@ 5.3.3. Resource Utilization and Overhead Breakdown
These experiments suggest that while Fisher computation contributes a consistent overhead during local training, communication latency remains the dominant factor in federated learning deployments (zhu2021delayed, ; le2024exploring, ). This indicates that curvature-based regularization is computationally feasible in practice (khan2025functional, ), with networking emerging as the primary bottleneck. Exploring lighter curvature proxies (wang2024zeroth, ) and adaptive scheduling (zhang2023timelyfl, ) remains a valuable direction for further improving scalability.
5.4. Sensitivity Analysis
  1. (1)
    Effect of Ordered Sampling on FL Performance
  2. (2)
  3. (3)
    Impact of Random Client Participation on Fairness Across Models
5.4.1. Effect of Ordered Sampling on FL Performance:
Human-sensing data often contains strong temporal dependencies (lymberopoulos2009methodology, ; li2025survey, ), making it important to preserve data order during local training (liu2024personalized, ; mashhadi2022auditing, ). It also mimics real-world FL deployments (wu2024spatio, ). To reflect this, we evaluate FL performance using an ordered 80:20 train-test split without shuffling. As shown in Table 8, CurvFed achieves the best fairness–utility tradeoff, reducing the EO gap and achieving the highest F1 and FATE scores, demonstrating its effectiveness in temporally structured FL settings.
Dataset Model F1 Score EO Gap FATE Score $\times 10^{-1}$
WIDAR FedAvg 0.8570 0.4629 0.0000
FedSRCVaR 0.7676 0.4002 0.311
KD-FedAvg 0.8135 0.4069 0.701
CurvFed 0.7488 0.3452 1.280
Stress Sensing FedAvg 0.7482 0.4000 0.0000
FedSRCVaR 0.6989 0.3633 0.259
KD-FedAvg 0.7326 0.3633 0.709
CurvFed 0.7693 0.2938 2.935
Table 8. Comparison of model performance under ordered 80:20 splits to capture temporal dependencies
In real-world edge deployments, clients often operate under resource constraints, leading to limited local data and sporadic client availability (li2021fedmask, ). To simulate this, we evaluate CurvFed under (i) varying local training data availability (20%, 50%, 90%), and (ii) fluctuating client participation per round (3, 5, 7 clients), with a fixed 20% test set.
(a) WIDAR Dataset
(b) Stress sensing Dataset
Figure 3. Impact of limited local data availability (20%, 50%, 90%) on fairness and utility in FL
(a) WIDAR Dataset
(b) Stress sensing Dataset
Figure 4. Impact of client participation subsampling (3 or 5 or 7 clients per round) on fairness–utility trade-off
5.4.3. Impact of Random Client Participation on Fairness:
In real-world FL deployments, especially in human-sensing applications, client availability is often inconsistent due to network instability and device constraints (lim2020federated, ). Unlike idealized settings with fixed clients per round, edge devices may randomly join or drop out during training. We simulate this, where clients were independently sampled using Bernoulli sampling with each client having a random probability (ranging from $0.1$ to $0.9$) of getting selected, introducing realistic variations in both the number and identity of participating clients.
Figure 5. Experiment on real-world FL evaluation simulating an inconsistent number of clients in each round.
5.5. Ablation Study
To validate the design choices of our proposed CurvFed framework, we conducted an extensive ablation study, detailed in Appendix. This study systematically isolates and evaluates the independent contributions of three key components: (1) the sharpness-aware aggregation strategy, (2) the sharpness-aware client-side training with fairness constraints, and (3) the use of the top eigenvalue of FIM penalty as a regularization term.
5.5.1. Independent Effect of Sharpness-Aware Aggregation:
We evaluated three aggregation strategies: based on classification loss, based on the top eigenvalue of the Fisher Information Matrix (FIM), and one that combines both via soft weighting—without applying FIM regularization during local client training. The combined approach consistently improved both fairness and accuracy across datasets, with an average EO Gap reduction of 6.78% compared to FedAvg, highlighting the standalone effectiveness of sharpness-aware aggregation.
5.5.2. Independent Effect of Sharpness-Aware Client-Side Training:
We applied the FIM top eigenvalue penalty during local client training while using a benign server-side aggregation strategy. This alone improved fairness, reducing the EO gap by an average of 6.60% compared to benign Fedavg across all three datasets, confirming the standalone benefit of sharpness-aware local training.
5.5.3. Effect of Top eigenvalue of FIM and SAM
We examined the impact of combining the client-side FIM top eigenvalue penalty with sharpness-aware aggregation and SAM. The results show:
  • The FIM penalty improves fairness but may slightly reduce accuracy.
  • SAM enhances model accuracy but slightly reduces fairness.
  • Combining the FIM penalty with sharpness-aware aggregation (without SAM) yields a better fairness–accuracy trade-off, though at a slight cost to accuracy.
5.5.4. Effectiveness of complete CurvFed Integration:
The full CurvFed configuration—integrating SAM, FIM top eigenvalue regularization, and sharpness-aware aggregation—achieves the best overall trade-off between fairness and efficacy. It consistently outperforms other variations across all datasets in terms of both EO Gap and F1 score, supporting the effectiveness of the complete approach described in Algorithm 1.
6. Empirical Justification
Figure 6. FATE score, F1 score, and loss landscape across communication rounds for CurvFed on Stress Sensing and WIDAR datasets
(a) Stress-sensing Dataset
(b) WIDAR Dataset
Figure 7. Comparison of Mean Person-wise top eigenvalues for Hessian and Fisher matrices across Rounds
Dataset Model $\Delta\mathrm{\lambda(\textbf{F}})$ $\Delta\mathrm{\lambda(\textbf{H}})$ FATE Score $\times 10^{-1}$
WIDAR Fedavg 0.009 0.061 0.0000
FedSAM 0.009 0.0365 -0.56
FedSWA 0.002 0.019 0.33
FedSRCVaR 0.0015 0.0310 0.160
KD-Fedavg 0.0022 0.0377 0.122
CurvFed 0.0007 0.0017 1.233
Stress Sensing Fedavg 0.03 0.0330 0.0000
FedSAM 0.019 0.0347 0.812
FedSWA 0.020 0.0278 0.38
FedSRCVaR 0.019 0.1446 0.319
KD-Fedavg 0.021 0.6089 1.645
CurvFed 0.004 0.0097 2.559
Table 9. Disparity based on ‘Sex’ Across Datasets for Different Models
Figure 8. Top eigenvalue of Fisher matrix by Sex and Hand across different clients and rounds in Stress Sensing Dataset.
(a) Client-wise comparison of maximum Fisher eigenvalue $\lambda(F_{\text{client}})$ and classification loss across different models
7. Broader Impact and Limitation
  • Achieving fairness in FL systems requires attaining a balance between fairness and model performance (gu2022privacy, ). The quest to ensure fairness for individual clients often results in reduced performance across the board, an issue highlighted in literature (ezzeldin2023fairfed, ; zhang2020fairfl, ; al2014crowd, ; wang2019age, ). Although CurvFed demonstrates significant effectiveness in achieving this balance, as discussed in Section 5.2, RQ-3, it exhibits minor performance degradation in the WIDAR and Stress sensing dataset for single-person clients. This could be attributed to our aggregation scheme, outlined in Section 4.1 Equation 12, which equally weighs low error rates and lower maximal uncertainty, thus limiting the ability to balance accuracy and fairness when dealing with single-person clients. The issue could be resolved by introducing a hyperparameter as the weighting factor.
  • Recent studies have explored the relationship between fairness and privacy in machine learning, often treating them as competing metrics (tran2023interplay, ; zhang2024unraveling, ). While this work has focused on fairness in FL, extending the investigation to explore the interplay between fairness and privacy in FL would be an important direction for future research.
  • A limitation of our study is the experimental scale, constrained by the lack of large-scale human-sensing datasets with sensitive attributes. While prior (non-human-centric) FL fairness works often evaluate on a larger number of clients, such datasets remain scarce in the human sensing domain (stateofalgobias, ). Even recent large-scale datasets like GLOBEM (xu2023globem, ) do not release demographic or other sensitive information needed for fairness evaluation. This gap has also been discussed in recent survey studies within the ubiquitous sensing community (yfantidou2023beyond, ; uncoveringbias, ).
  • Like many Rawlsian or min–max fairness approaches (lahoti2020fairness, ; hashimoto2018fairness, ), CurvFed promotes equity by aligning loss landscape curvatures across clients, which prioritizes worst-case (or underrepresented) groups. This strategy can occasionally come at the expense of performance in majority groups, a trade-off that has also been reported in other fairness-driven FL methods (mohri2019agnostic, ; papadaki2022minimax, ; du2021fairness, ). While this reflects a broader property of parity-based optimization (foulds2020parity, ; padh2021addressing, ) rather than a limitation of CurvFed itself, it highlights an important avenue for future research: designing FWD mechanisms that uplift the worst-off groups without diminishing performance for others, thereby moving toward Pareto-improving fairness (zanna2024enhancing, ; nagpal2025optimizing, ).
8. Conclusion
A.1. Dataset and Model Description
A.1.1. Dataset Description
WIDAR Dataset: The WIDAR3.0 dataset (zhang2021WIDAR3, ) contains WIFI channel state information (CSI) (kaurwireless, ) which represents the variation in the Wifi channel portrayed by commodity wifi devices (kaurwireless, ; zhang2021WIDAR3, ). The WIFI CSI contains fine-grained information about variations in the Wifi channel based on the Doppler phenomenon (zhang2021WIDAR3, ). These signals have been used by prior researchers for gesture detection (zhang2021WIDAR3, ), presence detection (sathya2023systematic, ), pose estimation (wang2018csi, ), human identification and activity detection (shahbazian2023human, ). We have used it as one of the HAR (Human Activity Recognition) datasets. The dataset contains data from 16 participants, 4 Females, and 12 Males, with five different orientations (labeled as 1 to 5). We removed a female participant from the dataset due to the absence of ground truth. The selected two gestures, ‘Slide’ and ‘Sweep,’ had data in all orientations for male and female participants in the presence of the wifi sensor. Since all participants had an equal number of samples in different orientations, i.e., orientations 1 to 5, we created a disparity in the dataset based on orientations by randomly sub-sampling 50% of the data belonging to orientations 4 and 5. This was done to create a training imbalance representative of real-life scenarios where some groups have more representation compared to others. We considered labels 1 to 3 as major orientation and the rest as minor orientation. For our work, we have trained a model to perform binary classification to detect Sweep vs. Slide gestures using WIFI-CSI data and fairness on the basis of ‘sex’ and ‘orientation’ as sensitive attributes.
HuGaDB Dataset: We have also utilized the HuGaDB (Human Gait Database) dataset (chereshnev2018hugadb, ) for HAR purposes, which comprises a diverse array of human activity recordings from 18 participants. The participant group includes 4 females and 14 males. In our work, we trained a binary classification model to distinguish between Standing and other activities from this dataset, with fairness considerations based on ‘sex’ as the sensitive attribute.
Stress sensing Dataset: Xiao et al. (xiao2024reading, ) collected this dataset. This dataset was obtained from 48 individuals, including 40 females and 8 males, who participated in four different stress-inducing tasks (Arithmetic, Bad Memory, Stress Videos, Pre-condition ), each associated with unique stressors. The dataset is not publicly available yet; Access to the dataset has been obtained through proper IRB approvals/extensions. Due to the continuation of data collection by the Xiao et al. (xiao2024reading, )’s research group, this paper’s reported dataset has more individuals than reported by the Xiao et al. (xiao2024reading, ) paper. The dataset has individuals wearing Empatica E4 (mccarthy2016validation, ) (collecting Stresssensing data) in their left and/or right hands and individuals’ sex information. Hence, this paper utilized ‘hand’ information and ‘sex’ as sensitive attributes for evaluation. Utilizing this dataset, we performed a binary stress detection classification task.
A.1.2. Client Distribution:
Distinct client distributions are established for each dataset in our study.
  • To effectively benchmark the performance of our proposed model and ensure clients with diverse participant groups, we deliberately structured the client distribution to include at least one participant from each sensitive attribute category. Table 10 provides a detailed overview of participant distribution across clients, categorizing them based on the sensitive attributes of males and females.
  • Given that our models do not depend on sensitive attribute information during training, we also demonstrate their effectiveness by emulating the approach in clients with either single and/or multiple participants. This intentional diversification, as depicted in Table 11, does not guarantee the presence of both sensitive attributes within each client.
Dataset Clients Male Female
Client 1 4 1
WIDAR Client 2 4 1
Client 3 4 1
Client 1 4 1
HugaDB Client 2 4 1
Client 3 4 1
Client 4 2 1
Client 1 2 4
Client 2 4 2
Client 3 1 6
Stress sensing Client 4 1 6
Client 5 2 3
Client 6 1 7
Client 7 3 6
Table 10. Random distribution of multi-person clients of WIDAR, HugaDB, and Stress sensing Datasets.
Dataset Clients Male Female
Client 1 1 1
Client 2 2 1
WIDAR Client 3 1 0
Client 4 2 0
Client 5 2 1
Client 6 1 0
Client 7 1 0
Client 8 2 0
Client 1 6 0
HugaDB Client 2 5 1
Client 3 0 2
Client 4 0 1
Client 1 1 0
Client 2 3 0
Client 3 6 10
Stress sensing Client 4 0 6
Client 5 3 7
Client 6 2 9
Client 7 0 1
Table 11. Single and multi-person clients distribution across WIDAR, HugaDB and Stress sensing Datasets
A.1.3. Model Description
WIDAR Model: The WIDAR Model architecture (adapted from (yang2023benchmark, )) is comprised of two fully connected layers within a sequential module, featuring an input size of 22 * 20 * 20. A dropout layer with a 40% dropout rate and a Rectified Linear Unit (ReLU) activation function follow the first linear layer, while the second linear layer outputs the classification result. The model reshapes the input tensor during the forward pass and incorporates dropout layers to mitigate over-fitting.
HugaDB Model: The HugaDB model architecture (adapted from (pirttikangas2006feature, )) is characterized by its complexity, featuring two sequential fully connected layers. The first layer consists of linear transformations with input and output sizes of 792 to 1024, followed by dropout with a 30% probability, Rectified Linear Unit (ReLU) activation, and Instance Normalization. Subsequently, another linear layer reduces the dimensionality to 512, followed by similar dropout, ReLU, and Instance Normalization. The second layer further reduces the dimensionality from 512 to 128, followed by a linear layer with an output size equal to the specified number of classes. The final output is obtained by applying the sigmoid function to the result. During the forward pass, the input tensor is processed through these layers, reshaped when necessary, and the output is returned after applying the sigmoid function.
Stress sensing Model: The Stress sensing model architecture (adapted from (sharma2022psychophysiological, )) comprises three sequential fully connected layers. The first layer takes input of size 34 and produces an output of size 64, incorporating batch normalization and Rectified Linear Unit (ReLU) activation. The second layer further processes the output, reducing it to size 128 with ReLU activation. The final layer transforms the output to the specified embedding size.
A.1.4. Hyperparameters
We have produced results from three different seed sets and presented one of them here. The complete set of results for all three seeds can be found in Table 12.
Dataset F1 Score Accuracy EO Gap Model Seed
0.8568 0.8501 0.4192 fedavg 123
0.8702 0.8654 0.3988 CurvFed 123
HugaDB 0.8496 0.8425 0.4369 fedavg 16
0.8856 0.8819 0.3952 CurvFed 16
0.8548 0.8476 0.4318 fedavg 12345
0.8896 0.8859 0.4012 CurvFed 12345
0.7466 0.7473 0.3316 fedavg 123
Stress sensing 0.7901 0.7868 0.2469 CurvFed 123
0.7383 0.7385 0.3086 fedavg 12345
0.7944 0.7912 0.2557 CurvFed 12345
0.7235 0.7253 0.3069 fedavg 2023
0.7814 0.7780 0.2522 CurvFed 2023
0.8645 0.8643 0.4199 fedavg 123
WIDAR 0.7818 0.7906 0.2983 CurvFed 123
0.8645 0.8643 0.4136 fedavg 12345
0.7897 0.7960 0.3086 CurvFed 12345
0.8507 0.8602 0.4264 fedavg 2023
0.8100 0.8123 0.3534 CurvFed 2023
Table 12. Performance metrics (F1, Accuracy, EO Gap) for FL models with different seeds across datasets
The approach uses several hyperparameters for Benign and CurvFed model training. Table 13 and 14 shows the hyperparameter values for benign and CurvFed model training, respectively. These values correspond to the best results obtained using a grid search. Within the context of the CurvFed model, the SWA Learning Rate specifically denotes the learning rate at which the SWA model is introduced into the training process. The SWA Start Round is indicative of the round at which the SWA mechanism is initiated. Additionally, the cycle parameter signifies that post the initiation of SWA, the global model and learning rate undergo updates using SWA in a cyclic manner, with each cycle representing a distinct iteration of this process. The term $\epsilon$ pertains to the weighting value, which ensures the maintenance of the relative ranks of each client during the aggregation process. Opting for a smaller value of $\epsilon$ proved beneficial for better ranking precision.
Hyperparameter WIDAR Stress sensing HugaDB
Learning Rate 0.1 0.1 0.001
Epochs 3 3 3
Total Round 80 80 100
Table 13. Hyperparameters for benign Fedavg model
Hyperparameter WIDAR Stress sensing HugaDB
Learning Rate 0.01 0.01 0.001
SWA Learning Rate 0.1 0.001 0.001
SWA start round 16 16 16
cycle 5 5 3
Epochs 3 3 3
alpha 0.92 0.92 0.92
$\epsilon$ 0.005 0.005 0.005
Total Rounds 80 80 80
Table 14. Hyperparameters for proposed CurvFed model
Appendix B Experiment Details
B.1. Details of ablation study
We conducted a comprehensive ablation study to meticulously guide the choices in our proposed approach. The study consistently demonstrates the effectiveness of our model across all datasets.
Dataset Client-Side Aggregation F1 Score EO Gap FATE Score $\times 10^{-1}$
Training Strategy
Benign Fedavg 0.8507 0.4263 0
SAM Fedavg 0.8005 0.4251 -0.560
WIDAR SAM Fedswa 0.8405 0.4071 0.331
SAM+ $\lambda(F_{client})$ (Eq11) Fedswa 0.8459 0.4058 0.424
$\lambda(F_{client})$ (Eq11) Fedavg 0.8253 0.3994 0.332
$\lambda(F_{client})$ (Eq11) $W^{r}$(Eq12) 0.8344 0.4135 0.109
[…]
$\lambda(F_{client})$ (Eq11) Sharpness Aware aggregation 0.8409 0.3982 0.545
SAM Fedswa+S(L) (Eq12) 0.8737 0.4199 0.420
SAM Fedswa+S(T) (Eq12) 0.8590 0.4084 0.518
SAM Sharpness Aware aggregation 0.8668 0.4161 0.429
[…]
Benign Fedavg 0.7253 0.3069 0
SAM Fedavg 0.7758 0.3034 0.812
Stress sensing SAM Fedswa 0.7406 0.3015 0.384
[…]
$\lambda(F_{client})$ (Eq11) Fedavg 0.7372 0.2998 0.395
$\lambda(F_{client})$ (Eq11) $W^{r}$(Eq12) 0.7345 0.2839 0.874
[…]
$\lambda(F_{client})$ (Eq11) Sharpness Aware aggregation 0.7077 0.2380 1.999
SAM Fedswa+S(L) (Eq12) 0.8091 0.2680 2.240
SAM Fedswa+S(T) (Eq12) 0.7995 0.2627 2.46
SAM Sharpness Aware aggregation 0.8224 0.2710 2.507
[…]
Benign Fedavg 0.8496 0.4369 0
SAM Fedavg 0.8555 0.4255 0.329
HugaDB SAM Fedswa 0.8418 0.4242 0.192
[…]
$\lambda(F_{client})$ (Eq11) Fedavg 0.8567 0.4142 0.602
$\lambda(F_{client})$ (Eq11) $W^{r}$(Eq12) 0.8574 0.4083 0.746
[…]
$\lambda(F_{client})$ (Eq11) Sharpness Aware aggregation 0.8510 0.4333 0.0982
SAM Fedswa+ S(L) (Eq12) 0.8934 0.4130 1.059
SAM Fedswa+S(T) (Eq12) 0.8888 0.4142 0.878
SAM Sharpness Aware aggregation 0.8911 0.4095 1.11
[…]
Table 15. Performance metrics on the different datasets for different FL strategies and model configurations
B.1.1. The Independent Effect of Sharpness-aware Aggregation and its Ablation Analysis
This section evaluates the effect of Sharpness-aware aggregation on enhancing FWD independently.
Previous studies have explored benign client-side training, using various weighted aggregation schemes to promote fairness (ezzeldin2023fairfed, ). CurvFed, for instance, aligns with Pessimistic Weighted Aggregation (P-W), a method that excludes highly biased (based on known bias-creating factors) models during aggregation. However, CurvFed takes a different approach, assigning lower weights to highly biased (based on loss-landscape sharpness; without the knowledge of the bias-creating factors) and low-efficacy local client-side models, thereby promoting fairness without completely excluding them.
To understand the benefits of CurvFed’s Sharpness-aware aggregation strategy independently (without client-side local training fairness constraints) in promoting clients with both (1) low error rates and (2) lower maximal uncertainty we evaluated three following variations, while all of them maintaining benign local client-side training without any fairness objectives.
  • CurvFed’s aggregation strategy as outlined in Equation 12 of the main paper while maintaining benign local client-side training without any fairness objectives.
  • CurvFed’s aggregation strategy modification, where just selecting clients with lower maximal uncertainty in prediction (Based on $T$ in Equation 12), while maintaining benign local client-side training without any fairness objectives.
  • CurvFed’s aggregation strategy modification, where just selecting clients with lower error rate (Based on $L$ in Equation 3), while maintaining benign local client-side training without any fairness objectives.
The evaluation results are presented in rows highlighted in ‘blue’ in Table 15. First, these results consistently demonstrate that across all datasets, CurvFed’s aggregation strategy, as outlined in Equations 2 and 3, outperforms others. Meaning this evaluation confirms that integrating both classification and top eigenvalue loss and favoring clients with lower loss (as outlined in Equations 2 and 3) enhances the balance between efficacy and fairness.
Furthermore, the fact that fairer models are achieved even without local client-side fairness constraints in these evaluations, compared to the benign baselines presented in rows highlighted in ‘red’ in Table 15, establishes that CurvFed’s Sharpness-aware aggregation independently promotes FWD.
B.1.2. The Independent Effect of Sharpness-aware Client-side Local Training
This section evaluates the independent effect of Sharpness-aware client-side local training as outlined in Equation 1 to enhance FWD. To conduct this evaluation, we implemented a variant where client-side local training incorporates fairness constraints based on Equation 1, while employing a benign aggregation scheme where all weights in Equation 2 are set to ‘1.’ To ensure a fair comparison, the SWA aggregation strategy, outlined in Algorithm 1, remains employed in this approach. The evaluation results are showcased in rows highlighted in ‘yellow’ in Table 15, indicating its superior enhancement of fairness compared to the benign approaches depicted in ‘red’ highlighted rows across all datasets. This underscores the independent influence of client-side sharpness-aware local training in enhancing FWD.
B.1.3. Independent Effect of FIM penalty term in Client-side training and Aggregation
In this section, we explore the independent effect of FIM penalty in the client-side training process and its interaction with various aggregation strategies. In our experiments, marked in orange, we observed that incorporating top eigenvalue of FIM either in the client-side training or aggregation, or having both without sharpness-aware strategies like SAM and SWA, leads to a fairer model, as indicated by a reduced EO Gap. However, this improvement in fairness comes at the cost of model accuracy, which tends to decrease.
The next observation shows that SAM significantly boosts the accuracy of the model, but this improvement comes at the cost of fairness, as reflected by a greater EO Gap. This suggests that SAM enhances model performance in terms of accuracy but introduces a trade-off in fairness.
For example, in the WIDAR dataset, applying the penalty term FIM penalty in client-side training while keeping the aggregation strategy benign reduces the EO Gap from 0.4263 (Benign) to 0.4186, reflecting a fairness improvement. However, this comes with a decline in F1 score, dropping from 0.8507 to 0.8394. Replacing the benign aggregation with a weighted strategy ($W^{r}$) further improves fairness, reducing the EO Gap to 0.4135, though the F1 score slightly decreases again to 0.8359. In contrast, using SAM in client-side training increases the F1 score to 0.8587, indicating enhanced model performance. Yet, this setting shows limited improvement in fairness, with the EO Gap remaining relatively high at 0.4212. This demonstrates that while SAM is effective for boosting accuracy, it does little to mitigate fairness issues. Interestingly, a more balanced configuration emerges when FIM penalty is applied without SAM in training and paired with sharpness aware aggregation. This setup achieves an F1 score of 0.8574 while also keeping the EO Gap at 0.4148, offering a better trade-off between fairness and accuracy. In the Stress Sensing dataset, applying the regularization term FIM penalty in client-side training while using FedAvg aggregation leads to a significant reduction in EO Gap, from 0.3069 (Benign) to 0.2539, indicating improved fairness. However, this results in a drop in F1 score from 0.7253 to 0.7034, reflecting a loss in model performance. On the other hand, SAM increases the F1 score to 0.7758, the highest among all settings, but only slightly improves fairness (EO Gap = 0.3034), showing again that SAM prioritizes accuracy over fairness. In the HugaDB dataset, applying FIM penalty with FedAvg yields the lowest EO Gap of 0.3928, improving fairness by over 10% compared to the baseline (0.4369), and even increases the F1 score to 0.8634, surpassing the benign setup. However, using SAM alone pushes the F1 score higher to 0.8789, but fairness drops slightly with EO Gap rising to 0.4000.
These findings emphasize the importance of incorporating our top eigenvalue of FIM regularization term in local training and aggregation strategies in server with sharpness-aware techniques like SAM or SWA. By doing so, we can strike a better balance between fairness and accuracy, ensuring the model performs well on both metrics.
B.1.4. Importance of Integrating Sharpness-Aware Client-side Training and Aggregation Strategy for FWD
The evaluations discussed above have demonstrated that both the Sharpness-Aware Aggregation Strategy, outlined in Equations 2 and 3, and the Sharpness-Aware client-side local training, depicted by Equation 1 in the main paper, independently contribute to promoting fairness. However, these individual strategies exhibit less effectiveness (fairness and efficacy balance) compared to the presented CurvFed approach, which incorporates both strategies, as indicated by the ‘green’ highlighted rows across all datasets in Table 15. This underscores the effectiveness of Algorithm 1 in attaining FWD.

### v7

@@ B.0.1. Resource Utilization and Overhead Breakdown
These experiments suggest that while Fisher computation contributes a consistent overhead during local training, communication latency remains the dominant factor in federated learning deployments (Zhu et al., 2021; Le et al., 2024). This indicates that curvature-based regularization is computationally feasible in practice (Khan et al., 2025), with networking emerging as the primary bottleneck. Exploring lighter curvature proxies (Wang et al., 2024) and adaptive scheduling (Zhang et al., 2023a) remains a valuable direction for further improving scalability.
Appendix C Experiment Details
C.1. Ablation Study on Additional Datasets
We conducted a comprehensive ablation study to meticulously guide the choices in our proposed approach. The study consistently demonstrates the effectiveness of our model across all datasets. This section presents the results for the Stress sensing, HugaDB and the WIDAR dataset in Table 18. The results for PERCEPT-R were reported in the main text under Section 6.3
# Dataset Client-Side Training Aggregation Strategy F1 Score ($\uparrow$) EO Gap ($\downarrow$) FATE Score $\times 10^{-1}$ ($\uparrow$)
1 WIDAR Benign FedAvg 0.8507 0.4263 0
2 SAM FedAvg 0.8005 0.4251 -0.560
3 SAM FedSWA 0.8405 0.4071 0.331
4 SAM + Eq. 11 FedSWA 0.8459 0.4058 0.424
5 (Eq11) FedAvg 0.8253 0.3994 0.332
6 (Eq11) $W^{r}$ (Eq12) 0.8344 0.4135 0.109
7 SAM + Eq. 11 $W^{r}$ (Eq12) 0.8707 0.4199 0.386
8 (Eq11) Sharpness-Aware Aggregation 0.8409 0.3982 0.545
9 SAM FedSWA + S(L) (Eq12* partial) 0.8737 0.4199 0.420
10 SAM FedSWA + S(T) (Eq12* partial) 0.8590 0.4084 0.518
11 SAM Sharpness-Aware Aggregation 0.8668 0.4161 0.429
12 SAM + Eq. 11 Sharpness-Aware Aggregation 0.8100 0.3533 1.233
13 Stress Sensing Benign FedAvg 0.7253 0.3069 0
14 SAM FedAvg 0.7758 0.3034 0.812
15 SAM FedSWA 0.7406 0.3015 0.384
16 SAM + Eq. 11 FedSWA 0.7346 0.2768 1.105
17 (Eq. 11) FedAvg 0.7372 0.2998 0.395
18 (Eq. 11) $W^{r}$ (Eq12) 0.7345 0.2839 0.874
19 SAM + Eq. 11 $W^{r}$ (Eq12) 0.8193 0.2927 1.756
20 (Eq. 11) Sharpness-Aware Aggregation 0.7077 0.2380 1.999
21 SAM FedSWA + S(L) (Eq12* partial) 0.8091 0.2680 2.240
22 SAM FedSWA + S(T) (Eq12* partial) 0.7995 0.2627 2.460
23 SAM Sharpness-Aware Aggregation 0.8224 0.2710 2.507
24 SAM + Eq. 11 Sharpness-Aware Aggregation 0.7814 0.2522 2.555
25 HugaDB Benign FedAvg 0.8496 0.4369 0
26 SAM FedAvg 0.8555 0.4255 0.329
27 SAM FedSWA 0.8418 0.4242 0.192
28 SAM + Eq. 11 FedSWA 0.8170 0.4142 0.135
29 (Eq. 11) FedAvg 0.8567 0.4142 0.602
30 (Eq. 11) $W^{r}$ (Eq12) 0.8574 0.4083 0.746
31 SAM + Eq. 11 $W^{r}$ (Eq12) 0.9016 0.4154 1.103
32 (Eq. 11) Sharpness-Aware Aggregation 0.8510 0.4333 0.098
33 SAM FedSWA + S(L) (Eq12* partial) 0.8934 0.4130 1.059
34 SAM FedSWA + S(T) (Eq12* partial) 0.8888 0.4142 0.878
35 SAM Sharpness-Aware Aggregation 0.8911 0.4095 1.110
36 SAM + Eq. 11 Sharpness-Aware Aggregation 0.8855 0.3952 1.376
Table 18. Ablation of CurvFed’s design choices on the WIDAR, Stress Sensing, and HugaDB datasets. Client-Side Training denotes the optimization or regularization applied locally at each client, while Aggregation Strategy specifies the server-side model aggregation rule. Unless explicitly stated, each configuration applies only the listed component, with no additional modifications. Metrics with $\uparrow$ indicate higher is better, while $\downarrow$ indicates lower is better.
We begin by establishing benign baselines using standard FedAvg aggregation without fairness-aware training (Rows 1, 13, and 25). Across datasets, these configurations achieve competitive F1 scores (0.8507 on WIDAR, 0.7253 on Stress Sensing, and 0.8496 on HugaDB) but are consistently associated with relatively high EO Gaps (0.4263, 0.3069, and 0.4369, respectively), highlighting the presence of fairness concerns under standard federated optimization.
Introducing SAM under FedAvg aggregation (Rows 2, 14, and 26) generally improves or maintains accuracy, most notably on Stress Sensing (F1 = 0.7758, Row 14), but yields only marginal changes in EO Gap. Replacing FedAvg with FedSWA while retaining SAM (Rows 3, 15, and 27) produces mixed effects, with modest reductions in EO Gap on WIDAR and Stress Sensing, but limited improvements on HugaDB, indicating that smoother aggregation alone is insufficient to consistently address bias.
We next examine the effect of enforcing curvature-aware local training without SAM. Applying the FIM-based regularization during client-side training with FedAvg aggregation (Rows 5, 17, and 29) consistently lowers EO Gap compared to the benign baselines, while also reducing F1 scores. Replacing FedAvg with curvature-aware weighting $W^{r}$ (Rows 6, 18, and 30) further modifies this trade-off, sometimes improving EO Gap (e.g., Row 18 on Stress Sensing) while slightly increasing F1 score relative to FedAvg-based curvature regularization. These results indicate that restricting local optimization to flatter regions promotes fairness, albeit at the cost of predictive performance when applied in isolation.
Combining SAM with curvature-aware local training and weighted aggregation (Rows 7, 19, and 31) shifts the balance toward higher accuracy. In particular, these configurations achieve some of the highest F1 scores across datasets, including the maximum F1 on HugaDB (0.9016, Row 31). However, EO Gaps remain relatively high in these settings, suggesting that SAM-driven optimization can counteract some fairness gains introduced by curvature regularization.
We then analyze aggregation-centric interventions. Applying sharpness-aware aggregation without SAM or curvature-aware local training (Rows 8, 20, and 32) yields noticeable reductions in EO Gap across datasets, most prominently on Stress Sensing (EO Gap = 0.2380, Row 20), but often with a reduction in F1 score. Partial sharpness-aware aggregation strategies under SAM, which prioritize either predictive uncertainty or classification loss (Rows 9–10, 21–22, and 33–34), achieve strong F1 scores—such as the highest F1 on WIDAR (0.8737, Row 9)—but do not consistently minimize EO Gap.
Employing the full sharpness-aware aggregation under SAM (Rows 11, 23, and 35) provides a more balanced outcome, achieving competitive F1 scores alongside moderate EO Gaps across all datasets. These results indicate that aggregation strategies that explicitly account for both loss and curvature information can mitigate bias more effectively than partial criteria.
Finally, the full CurvFed configuration—integrating curvature-aware client-side training with sharpness-aware aggregation—consistently achieves the most favorable fairness–utility trade-off. On WIDAR (Row 12), this configuration yields the lowest EO Gap (0.3533) and the highest FATE score (1.233). On Stress Sensing (Row 24), it achieves an EO Gap of 0.2522 with the highest FATE score (2.555). Similarly, on HugaDB (Row 36), it attains the lowest EO Gap (0.3952) and the highest FATE score (1.376), while maintaining a strong F1 score (0.8855). Although higher absolute F1 scores are observed in some partial configurations (e.g., Row 31), the full CurvFed model consistently offers the most balanced performance across datasets.